OpenStudy (katielong):

would (3sqrt3)squared be 3x3 which would be 9? i will put the equation on as an answer below. thanks =)

OpenStudy (anonymous):

no, 3^2*3=27

OpenStudy (katielong):

\[\left( 3\sqrt{3} \right)^{2}\]

OpenStudy (lgbasallote):

\[(3\sqrt 3)^2 = (3^2) \times (\sqrt 3^2) = 9 \times 3\]

OpenStudy (anonymous):

You have to square both numbers inside your parenthesis. As stated more quickly above me. :-)

OpenStudy (katielong):

arite, cheers everyone!! =)